OpenStudy (anonymous):

DIFFERENTIAL EQUATION find y given the differential equation (x+1)dy/dx = y with condition y(4)=10

zepdrix (zepdrix):

Hmm this appears to be separable. So let's move some things around.

zepdrix (zepdrix):

There is a process that allows us to write the dx differential on the other side. For our purposes, we can simply think of this as multiplication, even though that's not exactly accurate. :)\[\large (x+1)dy=ydx\]

zepdrix (zepdrix):

Now we'll move the y term over with the dy, and the x term over to the side with the dx.\[\large \frac{dy}{y}=\frac{dx}{x+1}\]

zepdrix (zepdrix):

And now we integrate! Yay!

zepdrix (zepdrix):

\[\large \int\limits \frac{dy}{y}=\int\limits \frac{dx}{x+1}\]

zepdrix (zepdrix):

\[\large \ln y = \ln(x+1)+c\]Confused about anything yet? Understand why we only write a +C on one side?

OpenStudy (anonymous):

I approached the question wrong before

OpenStudy (anonymous):

+C is what i need to solve right?

OpenStudy (anonymous):

the initial condition

zepdrix (zepdrix):

Yes, the initial condition allows you to find your C :)

zepdrix (zepdrix):

We don't want to find this C though. We want to manipulate the problem a little more BEFORE we solve for C.

zepdrix (zepdrix):

If we exponentiate both sides (Write them with a base of e) we get,\[\large e^{\ln y} = e^{\ln(x+1)+c}\]

zepdrix (zepdrix):

the exponential and logarithm are INVERSE operations of one another, so on the left, they'll cancel out nicely.\[\large y = e^{\ln(x+1)+c}\]

OpenStudy (anonymous):

then its y=(x+1)+c right?

zepdrix (zepdrix):

No. the right side you can't pull that little tricky I'm afraid, because of the +c attached to the log.

OpenStudy (anonymous):

ok thx for pointing thst out

zepdrix (zepdrix):

We have to do a little something involving exponents before we can do that little magic trick.

zepdrix (zepdrix):

Recall this exponential law?\[\huge a^{b+c} \quad = \quad a^b\cdot a^c\]

OpenStudy (anonymous):

im guessin we do e^(ln(x+1) + e^c or am i doing it wrong

OpenStudy (anonymous):

e^(ln(x+1) * e^c

zepdrix (zepdrix):

Applying this law lets us rewrite the right side,\[\huge y = e^{\ln(x+1)}\cdot e^c\]You're on the right track, just be careful with the operation you put when you brought it down. Ah yes good good c:

zepdrix (zepdrix):

NOW we're allowed to pull that little inverse tricky on the first part.

OpenStudy (anonymous):

e^c = c right

OpenStudy (anonymous):

since its an arbitrary constant?

zepdrix (zepdrix):

Yes, very good. e^c=C Given that C must be larger than 0.

OpenStudy (anonymous):

i get c = 5

OpenStudy (anonymous):

but the answer says y=2(x+1)

zepdrix (zepdrix):

Yah you should get 2, how did you get 5? I think you made a silly mistake maybe.

zepdrix (zepdrix):

\[\large y=Ce^{\ln(x+1)} \quad \rightarrow \quad y=C(x+1)\] \[\large y(4)=10 \quad \rightarrow \quad 10=C(4+1)\]

OpenStudy (anonymous):

its multiplied by c

OpenStudy (anonymous):

how come?

zepdrix (zepdrix):

Because of the Exponential Law that we applied, we wrote it as multiplication of bases. Remember the e term?

OpenStudy (anonymous):

oh i see

zepdrix (zepdrix):

You have to be careful with problems like this, because the constant can change from addition to multiplication, depending on what you do to the problem.

OpenStudy (anonymous):

i always make mistakes like that

zepdrix (zepdrix):

Heh :3 practice!

OpenStudy (anonymous):

I got it thx alot

zepdrix (zepdrix):

np.